Can You Help Me Solve This Expression? 14x-11x-6x
Answer by jim_thompson5910(35256)

These are all like terms because they all have an 'x' in them and nothing else. So we just combine the numbers in front and ignore the x terms. We get: 14-11-6=3-6=-3. After that, stick an 'x' to the end of it.


So the final answer is -3x